OpenStudy (anonymous):

John wants to rent a tent to go camping. The cost to rent a tent from Woodland Outfitters is a flat rate of $10.00 plus $0.50 for each night the tent is rented. Part a. Let x represent the number of nights John keeps the tent and write a cost function, c(x), for the cost to rent a tent from Woodland Outfitters. Part b. How much will John owe if he uses the tent for five nights?

OpenStudy (anonymous):

c(x)=10+(0.50)x c(5)=10+(0.50)(5) =12.50

OpenStudy (anonymous):

Thank you so much! :)